# Quantitative approaches for multi-scale structural analysis with atomic resolution electron microscopy

### v1

@@ III Local wave fitting for lattice parameter measurement
16 While damping information in Fourier space is useful for revealing superlattices associated with discrete broken symmetries, in many systems, continuous deformations which locally alter lattice parameters are key to understanding functional properties. The combination of high resolution and high precision make STEM a powerful technique for measuring such spatial variations in inter-planar spacings and angles. Often parameterized as strains - relative measures of deformation compared to a reference region or nominal value - an understanding of these variations on the nanometer- to micron-scale is frequently a primary goal of STEM analysis. Applications include identifying defects, checking for relaxation across interfaces, differentiating domains associated with twinning or ferroelasticity, identifying particular components of nanostructures, and device metrology [38, 39, 40, 41, 42, 43].
@@
17 In an atomic resolution STEM image, local strain can be measured simply by calculating the distances between neighboring atomic columns and comparing to a known value or reference region [7]. Mapping strain across a full image in this way, however, depends on precise identification of the positions of all of the columns in the image, involving slow, resource intensive fitting of each column, such that success depends strongly on the resolution and signal-to-noise ratio of the image. More commonly, strains are measured using GPA [20, 8] or phase-lock in approaches [44, 45, 33], which Fourier filter a single peak to select a periodicity of interest, calculate a geometric phase that captures the deviations from this periodicity, and measure strains from its gradient.
@@
18 By eliminating the need for precisely fitting the positions of each atomic column, these approaches allow analysis to be easily scaled to larger areas, but also have their own limitations. One is that the quantification is relative to a reference periodicity, complicating absolute measurement of inter-planar spacings. Another is the sensitivity of these measurements to noise: if the periodicity being studied has a low signal-to-noise ratio (SNR) - for instance because its amplitude is suppressed from disorder - artifacts which resemble phase slips and dislocations but have no physical origin will appear due to the nature of the image processing performed [32]. Artifacts can also emerge from changing contrast of atomic columns, for instance due to different compositions across heterostructure interfaces [46]. These issues are compounded by the lack of a means to readily measure the uncertainty or error associated with the measurement of the phase, meaning processing artifacts can easily be mistaken for real sample features. On the other hand, with recent 4D-STEM detector development, strain is now commonly measured with scanning nanodiffraction, which can experimentally be very convenient and requires no atomic resolution imaging at all [47]. Still, the resolution of this technique is limited by the requirement of a small convergence angle to avoid overlap of diffraction discs, which can be challenging to optimize especially in materials with large unit cells or highly localized strain variations. As the convergence angle must be selected prior to the measurement, fine features and the presence of defects which are clear in imaging may be missed.
@@
19 Here, we present a technique for quantifying lattice spacings and angles from atomic resolution images by locally fitting waves in real space to a Fourier filtered signal, developed for analyzing challenging nanostructured materials [22]. While the Fourier processing is similar to GPA, by fitting in real space this approach offers distinct advantages. For instance, absolute inter-planar spacings are measured rather than relative strains, eliminating the need for a pristine reference area for use as a reference. Additionally, the fitting variance can be used as a heuristic for the uncertainty associated with the measurement, providing the error bars missing from techniques like GPA as well as a means of identifying and excluding artifacts and unreliable measurements.
@@
20 Figure 5 shows the steps involved in the local wave fitting algorithm, using as an example an image of a partially relaxed thin film of Sr2RuO4 grown on a SrLaAlO4 substrate with an array of misfit dislocations along the film-substrate interface. Fourier filtering the in-plane (200) peak by weighting the Fourier transform of the STEM image with a narrow Gaussian distribution centered on the peak of interest, such that only information about that periodicity is retained, and taking the real part of the inverse transform back to real space results in sinusoidal fringes peaked on each (200) plane of the lattice. As detailed in Supplementary Note 1 [37], the spatial resolution or coarsening length associated with this Fourier filtering is inversely proportional to the width of the Gaussian distribution used. Increasing the filter size may also incorporate more noise, reducing the SNR and potentially introducing artifacts. Thus, an appropriately sized filter - wide enough that a sufficient spatial resolution is achieved to map the features of interest, but not so wide that more than one peak or excessive noise is included - is critical for meaningful analysis. As can be seen in Figure 5c which shows a portion of this fringe image around one of the misfit dislocations, variations in the periodicity of the wave reflect the different lattice spacings of the film and substrate, while the misfit dislocation is marked by additional half-planes and a bending of the surrounding fringes, matching the behavior of the atomic planes in the original image.
@@
21 The fringe image is then divided into small patches, typically around twice the fringe wavelength, so that variations in the wavelength and angle of the fringes are apparent. To avoid degradation in resolution, the sampling of the patches is made at a greater frequency than the patch size, such that they overlap (Fig. 5d). The amplitude, wavelength, and orientation of the planes in each patch are found by fitting with 2D sine waves, allowing the spatial variations in each of these parameters to be mapped. For instance, patches from the film have an increased wavelength compared to the average across the full field of view, which can be seen clearly when projected to one dimension (Fig. 5e,f), while fringes in the substrate are relatively compressed, reflecting its smaller in-plane lattice parameter (Fig. 5g,h).
@@
22 By fitting these waves in patches across the whole image, the spacing and orientation of the (200) planes can be mapped out from the fit parameters (Fig. 5i,j). The inter-planar spacing is directly measured in real space units, an important consideration when an understanding of absolute lattice parameters rather than relative strains is desired. In this case, the independent measurement of the film and substrate lattice parameters shows the expected uniform relaxation of the film suggested by the presence of the regularly spaced misfit dislocations, and an interesting visualization of the deformation surrounding the defects. By repeating the analysis with other peaks, other periodicities such as the out-of-plane spacing can be measured in the same way.
@@
Figure 6: Local wave fitting application and comparison with geometric phase analysis. a) A HAADF-STEM image of a superlattice of PbSe quantum dots. b) Fringes mapped from real part of Fourier transform of the $I_{2}$ peak marked in (d). c) Longitudinal strains (left and center) measured along two principal axes from wave fit d-spacings relative to the average spacings across the field of view, and the local lattice rotation (right). d) Fourier transform of the STEM image, showing azimuthally elongated Bragg peaks due to orientational disorder in the superlattice. Superlattice peaks can be seen very close to the central peak but are not used in this analysis. e) Phase associated with the Bragg $I_{2}$ peak marked in (d) extracted with geometric phase analysis. Large phase variations arise from real strains in the system as well as from artifacts in areas without crystalline structure. f) Strains mapped with geometric phase analysis, showing strong similarity to the wave fit result. g) Relative errors of $\epsilon_{xx}$ and $\epsilon_{yy}$ summed in quadrature, which peaks in gaps between quantum dots without crystaline order. h) Mask created by thresholding the fit error which segments areas with low fit certainty. i) Strains measured with wave fitting and masked to include only areas with low uncertainty.
@@
23 In cases where specifically a strain measurement is desired, the inter-planar spacings can be converted to relative strains simply by measuring the deviation in the measured wavelength relative to a chosen reference. To validate the approach, strain is measured in a superlattice of interconnected quantum dots [48, 22] with both the conventional GPA approach and the local wave fitting algorithm presented here (Fig. 6a). The strains measured with each technique are very comparable, revealing in both cases the variations in particle orientation and localized strains present at the crystalline bridges between some dots (Fig. 6c,f).
@@
24 This similarity is indeed unsurprising, as both measurements at a basic level are mapping phase variations in the same Fourier filtered signal. The main distinguishing feature of the different approaches is thus their interpretability. A key advantage of the local wave fitting approach and motivation for its use is the extraction of not only spacing and rotation of lattice planes, but also variances of the associated fits. These variances give valuable information about the error associated with lattice parameter or strain measurements, which is missing from conventional approaches like GPA and phase lock-in. This is particularly important when studying inhomogenous systems, such as nanostructures with distinct contrast in different components, twin domains with peaks at distinct positions in reciprocal space, and any images with significant local disorder or otherwise poor signal-to-noise ratios - all of which can give rise to artifacts in the form of non-physical phase variations. These uncertainties can be used to set independent, quantitative error bounds on all measured strains and parameters, allowing significant and insignificant variations to be differentiated and comparisons with other characterization techniques such as electron or X-ray diffraction to be better understood. For instance, the uncertainties can be used to understand the error in the strain measured over continuous and non-continuous necks between a pairs of quantum dots, shown in Supplementary Note 4 [37].
@@
25 For another example, GPA and local wave fitting both measure large apparent strains and discontinuities in the gaps between particles where there is no lattice present at all. These large “strains” and apparent defects emerge because when the amplitude of the Fourier filtered signal is low, the image’s phase information is dominated by noise, rather than real sample features. This noise causes the phase to vary rapidly, as can be seen in the image of the geometric phase extracted with GPA (Fig. 6e). Because these rapid variations can look quite similar to large strains associated with dislocations and out-of-phase boundaries, hasty analysis may confuse them for real features. The variance of the fits performed with wavefitting allow these unreliable areas to easily be identified and masked (Fig. 6h,i) in order to visualize only the meaningful strain information extracted from the image (Fig. 6j). While a similar mask could be created in this case by using the amplitude of the Fourier filtered signal, the rich, quantitative information provided by the fit uncertainty offers widely useful information for the interpretation of measured strains, a key advantage that other similar methods lack.
@@ IV Discussion
26 This work has presented approaches for reliable quantification of STEM images using a synergy of real and reciprocal space information to address two key structural characterization problems – the measurement of translational symmetry breaking distortions and local variations in lattice parameters. Taking advantage of the localization of information associated with symmetry breaking in reciprocal space, the Fourier damping algorithm facilitates the high-precision measurement of periodic lattice displacements, as well as a variety of other distortions including chemical and vacancy orderings. By enabling direct comparison between distorted and parent structures, the technique facilitates physically meaningful analysis. This ability to separate distortions of interest from lattice disorder and other coincident orderings helps to disentangle the structural complexities found in many functional materials. Similarly, the local wave fitting approach for lattice parameter quantification offers a convenient means for absolute measurement of inter-planar spacings and angles, and comparable strain measurements to common approaches such as GPA, while providing an enhanced understanding of the uncertainty associated with these measurements. These algorithms, along with a suite of other atomic resolution image processing routines including phase-lock in analysis, calculation of correlation functions, and other local crystallographic techniques have been implemented and made available to the community in the open-source kemstem Python package.
27 With the growing adoption of atomic resolution STEM characterization in the materials development cycle, the development of robust analysis techniques capable of reliably extracting the maximum available structural information from the resulting imaging data is essential. While advances in computing hardware and statistical analysis approaches enable the study of increasingly large and complex datasets, care must nonetheless be taken to ensure that the resulting insights are physically meaningful. The techniques presented here show how applying physically informed analysis to harness both the real and reciprocal space information encoded in atomic resolution STEM images can enable robust quantification of features of interest while transparently revealing the associated sources of error and levels of uncertainty. Continuing development in these areas will further improve the reliability and scalability of atomic resolution image analysis to introduce greater automation both for post-hoc analysis and real-time on-line processing, as well as adapting approaches to take full advantage of hardware developments in general purpose GPU computing.
@@ V Data Availability
The datasets analyzed in this study are available from the Platform for the Accelerated Realization, Analysis, and Discovery of Interface Materials (PARADIM) database at https://doi.org/10.34863/xxxxx[49].

### v2

@@ III Local wave fitting for lattice parameter measurement
While damping information in Fourier space is useful for revealing superlattices associated with discrete broken symmetries, in many systems, continuous deformations which locally alter lattice parameters are key to understanding functional properties. The combination of high resolution and high precision make STEM a powerful technique for measuring such spatial variations in inter-planar spacings and angles. Often parameterized as strains - relative measures of deformation compared to a reference region or nominal value - an understanding of these variations on the nanometer- to micron-scale is frequently a primary goal of STEM analysis. Applications include identifying defects, checking for relaxation across interfaces, differentiating domains associated with twinning or ferroelasticity, identifying particular components of nanostructures, and device metrology [38, 39, 40, 41, 42, 43].
@@
In an atomic resolution STEM image, local strain can be measured simply by calculating the distances between neighboring atomic columns and comparing to a known value or reference region [7]. Mapping strain across a full image in this way, however, depends on precise identification of the positions of all of the columns in the image, involving slow, resource intensive fitting of each column, such that success depends strongly on the resolution and signal-to-noise ratio of the image. More commonly, strains are measured using GPA [20, 8] or phase-lock in approaches [44, 45, 34], which Fourier filter a single peak to select a periodicity of interest, calculate a geometric phase that captures the deviations from this periodicity, and measure strains from its gradient.
@@
By eliminating the need for precisely fitting the positions of each atomic column, these approaches allow analysis to be easily scaled to larger areas, but also have their own limitations. One is that the quantification is relative to a reference periodicity, complicating absolute measurement of inter-planar spacings. Another is the sensitivity of these measurements to noise: if the periodicity being studied has a low signal-to-noise ratio (SNR) - for instance because its amplitude is suppressed from disorder - artifacts which resemble phase slips and dislocations but have no physical origin will appear due to the nature of the image processing performed [33]. Artifacts can also emerge from changing contrast of atomic columns, for instance due to different compositions across heterostructure interfaces [46]. These issues are compounded by the lack of a means to readily measure the uncertainty or error associated with the measurement of the phase, meaning processing artifacts can easily be mistaken for real sample features. On the other hand, with recent 4D-STEM detector development, strain is now commonly measured with scanning nanodiffraction, which can experimentally be very convenient and requires no atomic resolution imaging at all [47]. Still, the resolution of this technique is limited by the requirement of a small convergence angle to avoid overlap of diffraction discs, which can be challenging to optimize especially in materials with large unit cells or highly localized strain variations. As the convergence angle must be selected prior to the measurement, fine features and the presence of defects which are clear in imaging may be missed.
@@
Here, we present a technique for quantifying lattice spacings and angles from atomic resolution images by locally fitting waves in real space to a Fourier filtered signal, developed for analyzing challenging nanostructured materials [22]. While the Fourier processing is similar to GPA, by fitting in real space this approach offers distinct advantages. For instance, absolute inter-planar spacings are measured rather than relative strains, eliminating the need for a pristine reference area for use as a reference. Additionally, the fitting variance can be used as a heuristic for the uncertainty associated with the measurement, providing the error bars missing from techniques like GPA as well as a means of identifying and excluding artifacts and unreliable measurements.
@@
Figure 5 shows the steps involved in the local wave fitting algorithm, using as an example an image of a partially relaxed thin film of Sr2RuO4 grown on a SrLaAlO4 substrate with an array of misfit dislocations along the film-substrate interface. Fourier filtering the in-plane (200) peak by weighting the Fourier transform of the STEM image with a narrow Gaussian distribution centered on the peak of interest, such that only information about that periodicity is retained, and taking the real part of the inverse transform back to real space results in sinusoidal fringes peaked on each (200) plane of the lattice. As detailed in Supplementary Note I [30], the spatial resolution or coarsening length associated with this Fourier filtering is inversely proportional to the width of the Gaussian distribution used. Increasing the filter size may also incorporate more noise, reducing the SNR and potentially introducing artifacts. Thus, an appropriately sized filter - wide enough that a sufficient spatial resolution is achieved to map the features of interest, but not so wide that more than one peak or excessive noise is included - is critical for meaningful analysis. As can be seen in Figure 5c which shows a portion of this fringe image around one of the misfit dislocations, variations in the periodicity of the wave reflect the different lattice spacings of the film and substrate, while the misfit dislocation is marked by additional half-planes and a bending of the surrounding fringes, matching the behavior of the atomic planes in the original image.
@@
The fringe image is then divided into small patches, typically around twice the fringe wavelength, so that variations in the wavelength and angle of the fringes are apparent. To avoid degradation in resolution, the sampling of the patches is made at a greater frequency than the patch size, such that they overlap (Fig. 5d). The amplitude, wavelength, and orientation of the planes in each patch are found by fitting with 2D sine waves, allowing the spatial variations in each of these parameters to be mapped. For instance, patches from the film have an increased wavelength compared to the average across the full field of view, which can be seen clearly when projected to one dimension (Fig. 5e,f), while fringes in the substrate are relatively compressed, reflecting its smaller in-plane lattice parameter (Fig. 5g,h).
@@
By fitting these waves in patches across the whole image, the spacing and orientation of the (200) planes can be mapped out from the fit parameters (Fig. 5i,j). The inter-planar spacing is directly measured in real space units, an important consideration when an understanding of absolute lattice parameters rather than relative strains is desired. In this case, the independent measurement of the film and substrate lattice parameters shows the expected uniform relaxation of the film suggested by the presence of the regularly spaced misfit dislocations, and an interesting visualization of the deformation surrounding the defects. By repeating the analysis with other peaks, other periodicities such as the out-of-plane spacing can be measured in the same way.
@@
Figure 6: Local wave fitting application and comparison with geometric phase analysis. a) A HAADF-STEM image of a superlattice of PbSe quantum dots. b) Fringes mapped from real part of Fourier transform of the $I_{2}$ peak marked in (d). c) Longitudinal strains (left and center) measured along two principal axes from wave fit d-spacings relative to the average spacings across the field of view, and the local lattice rotation (right). d) Fourier transform of the STEM image, showing azimuthally elongated Bragg peaks due to orientational disorder in the superlattice. Superlattice peaks can be seen very close to the central peak but are not used in this analysis. e) Phase associated with the Bragg $I_{2}$ peak marked in (d) extracted with geometric phase analysis. Large phase variations arise from real strains in the system as well as from artifacts in areas without crystalline structure. f) Strains mapped with geometric phase analysis, showing strong similarity to the wave fit result. g) Relative errors of $\epsilon_{xx}$ and $\epsilon_{yy}$ summed in quadrature, which peaks in gaps between quantum dots without crystalline order. h) Mask created by thresholding the fit error which segments areas with low fit certainty. i) Strains measured with wave fitting and masked to include only areas with low uncertainty. Coarsening lengths associated with the geometric phase analysis (3.6 Å) and local wave fitting (5.7 Å) are marked by the diameters of white circles in the lower right of each panel, as discussed in Supplementary Notes 1 and 2 [30].
@@
In cases where specifically a strain measurement is desired, the inter-planar spacings can be converted to relative strains simply by measuring the deviation in the measured wavelength relative to a chosen reference. To validate the approach, strain is measured in a superlattice of interconnected quantum dots [48, 22] with both the conventional GPA approach and the local wave fitting algorithm presented here using identical 2D Gaussian filters (Fig. 6a). The strains measured with each technique are very comparable, revealing in both cases the variations in particle orientation and localized strains present at the crystalline bridges between some dots (Fig. 6c,f).
@@
This similarity is indeed unsurprising, as both measurements at a basic level are mapping phase variations in the same Fourier filtered signal. The main distinguishing feature of the different approaches is thus their interpretability. A key advantage of the local wave fitting approach and motivation for its use is the extraction of not only spacing and rotation of lattice planes, but also variances of the associated fits. These variances give valuable information about the error associated with lattice parameter or strain measurements, which is missing from conventional approaches like GPA and phase lock-in. This is particularly important when studying inhomogeneous systems, such as nanostructures with distinct contrast in different components, twin domains with peaks at distinct positions in reciprocal space, and any images with significant local disorder or otherwise poor signal-to-noise ratios - all of which can give rise to artifacts in the form of non-physical phase variations. These uncertainties can be used to set independent, quantitative error bounds on all measured strains and parameters, allowing significant and insignificant variations to be differentiated and comparisons with other characterization techniques such as electron or X-ray diffraction to be better understood. For instance, the uncertainties can be used to understand the error in the strain measured over continuous and non-continuous necks between a pairs of quantum dots, shown in Supplementary Note V [30].
@@
For another example, GPA and local wave fitting both measure large apparent strains and discontinuities in the gaps between particles where there is no lattice present at all. These large “strains” and apparent defects emerge because when the amplitude of the Fourier filtered signal is low, the image’s phase information is dominated by noise, rather than real sample features. This noise causes the phase to vary rapidly, as can be seen in the image of the geometric phase extracted with GPA (Fig. 6e). Because these rapid variations can look quite similar to large strains associated with dislocations and out-of-phase boundaries, hasty analysis may confuse them for real features. The variance of the fits performed with wave fitting allow these unreliable areas to easily be identified and masked (Fig. 6h,i) in order to visualize only the meaningful strain information extracted from the image (Fig. 6j). While a similar mask could be created in this case by using the amplitude of the Fourier filtered signal, the rich, quantitative information provided by the fit uncertainty offers widely useful information for the interpretation of measured strains, a key advantage that other similar methods lack.
@@ IV Discussion
This work has presented approaches for reliable quantification of STEM images using a synergy of real and reciprocal space information to address two key structural characterization problems – the measurement of translational symmetry breaking distortions and local variations in lattice parameters. Taking advantage of the localization of information associated with symmetry breaking in reciprocal space, the Fourier damping algorithm facilitates the high-precision measurement of periodic lattice displacements, as well as a variety of other distortions including chemical and vacancy orderings. By enabling direct comparison between distorted and parent structures, the technique facilitates physically meaningful analysis. This ability to separate distortions of interest from lattice disorder and other coincident orderings helps to disentangle the structural complexities found in many functional materials. Similarly, the local wave fitting approach for lattice parameter quantification offers a convenient means for absolute measurement of inter-planar spacings and angles, and comparable strain measurements to common approaches such as GPA, while providing an enhanced understanding of the uncertainty associated with these measurements. These algorithms, along with a suite of other atomic resolution image processing routines including phase-lock in analysis, calculation of correlation functions, and other local crystallographic techniques have been implemented and made available to the community in the open-source kemstem Python package.
With the growing adoption of atomic resolution STEM characterization in the materials development cycle, the development of robust analysis techniques capable of reliably extracting the maximum available structural information from the resulting imaging data is essential. While advances in computing hardware and statistical analysis approaches enable the study of increasingly large and complex datasets, care must nonetheless be taken to ensure that the resulting insights are physically meaningful. The techniques presented here show how applying physically informed analysis to harness both the real and reciprocal space information encoded in atomic resolution STEM images can enable robust quantification of features of interest while transparently revealing the associated sources of error and levels of uncertainty. Continuing development in these areas will further improve the reliability and scalability of atomic resolution image analysis to introduce greater automation both for post-hoc analysis and real-time on-line processing, as well as adapting approaches to take full advantage of hardware developments in general purpose GPU computing.
@@ V Data Availability
The datasets analyzed in this study are available from the Platform for the Accelerated Realization, Analysis, and Discovery of Interface Materials (PARADIM) database at https://doi.org/10.34863/mjbn-ym63 [49].